# Comparison of generalised additive models and neural networks in applications: A systematic review

### v1

@@ 1 Introduction
Over the past two decades, rapid advancements in machine learning and artificial intelligence have reshaped the abilities of predictive modelling. Among the most transformative of these developments are neural networks (Rosenblatt,, 1958; Rumelhart et al.,, 1986), which have demonstrated high predictive performance across a wide range of tasks, including image recognition, natural language processing, and forecasting (Abiodun et al.,, 2018; Liu et al.,, 2017). In many cases, the primary objective when employing neural networks is to maximise predictive accuracy, with less emphasis placed on model interpretability. These models are often regarded as “black-box” algorithms: although their interconnected and non-linear structure allows for high flexibility, it also makes for a model that is difficult to interpret. One of the key strengths of neural networks is that, even with a single hidden layer, they have been proven to be universal approximators (Cybenko,, 1989; Hornik et al.,, 1989). This property implies that neural networks can approximate any continuous, single-valued function, and hence could be strong predictors. However, these theories of universality may require the neural network size to grow to infinity. Recent studies have shown that deep neural networks of a fixed size, i.e., neural networks with more than one hidden layer, can also achieve universal approximation, often with a drastically reduced number of parameters (Lin et al.,, 2017; Mhaskar and Poggio,, 2016). This depth-induced parsimony improves computational and statistical efficiency.
@@
The majority of neural network research has been conducted outside of the field of statistics, resulting in a limited incorporation of statistically-based methodologies and a disparity in the terminology used (Hooker and Mentch,, 2021). Neural networks, particularly multilayer perceptrons, have emerged as strong alternatives to traditional statistical models, albeit they are often viewed as quite separate entities. However, a number of researchers have illustrated the similarities between neural networks and statistical models. Sarle, (1994) conducted a comprehensive comparison of various neural network models with traditional statistical approaches, showing that multilayer perceptrons can be viewed as generalisations of non-linear regression and discriminant models. Warner and Misra, (1996) highlight the parallels between neural networks and statistical regression models in terms of notation and implementation, showing that neural networks act as a type of nonparametric regression model. Cheng and Titterington, (1994) inform statistical readers about neural networks, highlighting similarities with statistical methodology. More recently, Ha and Burke, (2024) utilise deep neural networks within both a generalised non-linear modelling framework (extending classical generalised linear models) and a non-linear survival model (extending the Cox model). Synthesizing these views, it becomes evident that neural networks are a generalisation of statistical approaches whereby they automatically learn complex non-linear effects and interactions within their hidden layers; this is in contrast to the manual specification of interactions required in traditional statistical modelling, which frequently utilises simpler linear effects for practicality.
@@
Given previous work comparing the mathematical structure of statistical models and neural networks, the question then arises as to how these methods compare with each other in practical applications. Indeed, several authors have made strides at answering this over the years. Sargent, (2001) conducted a review and analysis of the literature comparing neural networks with linear, logistic, and Cox regression in the context of medical research. Karlaftis and Vlahogianni, (2011) provided a survey of applications in the transportation area, comparing neural networks with linear or logistic regression predominantly. Shin et al., (2021) carried out a systematic literature review of machine learning techniques (random forests, decision and regression trees, support vector machines, and neural networks) versus conventional statistical models (logistic, Cox, and Poisson regression) for predicting heart failure readmission and mortality. Cho et al., (2021) carried out a systematic review of the same aforementioned machine learning methods and conventional statistical models for the prediction of myocardial infarction readmission and mortality. Sun et al., (2022) conducted a systematic review and meta-analysis of machine learning techniques, including random forests, support vector machines, neural networks, and decision trees, against statistical methods, linear and Cox regression, for predicting heart failure events. More generally, Paliwal and Kumar, (2009) carried out a review of comparative studies across various application domains, evaluating multilayered feedforward neural networks against linear regression, logistic regression, and discriminant analysis. Their review critically assessed the literature based on several key criteria, including sample size, adherence to the assumptions of the techniques, the method by which results are validated, the metrics used for comparison, and whether statistically significant differences were observed in the outcomes.
@@
Despite the interest in comparing neural networks and statistical models, prior reviews have predominantly narrowed their view to particular subject areas, such as transportation and heart health, as mentioned above. Additionally, these reviews have focused on quite simple statistical models, such as linear regression, which inherently lack the capabilities to model non-linear effects that may be underlying in the data. Therefore, such comparisons are not fair representations of what is possible in more advanced non-linear statistical models. Specifically, Generalised Additive Models (GAMs) (Hastie and Tibshirani,, 1986) are the most popular extension to linear regression, replacing linear terms with arbitrary non-linear functions. The strength of GAMs is their ability to deal with non-linear relationships between predictor and response variables, while maintaining interpretability through their additive structure.
@@
To date, there has been no systematic review of the literature comparing GAMs with neural networks in a wide range of application areas, and, therefore, our current paper closes this gap. The purpose of the systematic review is to be a source of information on the types of studies that have carried out comparisons of these two model classes and the reported findings within these studies. To move beyond just a summary of these papers, we have further extracted and analysed the values of the performance metrics provided within the papers, along with characteristics that can potentially explain observed differences (e.g., application area, year of study, sample size, number of predictor variables, and neural network complexity).
@@ 2 Theoretical background
Before providing the details of our systematic review, we first provide some brief technical details on the models under consideration. Traditional statistical models and neural networks share the fundamental objective of modelling the relationship between a set of covariates and a response variable. Let $Y\in\mathbb{R}$ denote the response variable and $X=(X_{1},X_{2},\dots,X_{k})^{\top}\in\mathbb{R}^{k}$ the vector of covariates, where we restrict attention to tabular data. The general regression framework can be written as
@@
where GAMs utilise smooth functions $g_{j}(x_{j})$ to model non-linear relationships and
$\textbf{NN}(x_{1},x_{2},\dots,x_{k})$ is a fully connected neural network (NN).
@@ 2.1 Generalised additive models (GAMs)
GAMs were first introduced by Hastie and Tibshirani, (1986, 1990), extending linear regression by allowing non-linear relationships between the predictor variables $\mathbf{x}$, and the response $Y$. As seen in Equation 2.2, smooth functions $g(\cdot)$ replace the constant coefficients of linear regression, which increases flexibility. Smooth functions can be approximated as a linear combination of a set of known basis functions,
@@
where $P$ is the number of basis functions, controlling the maximum complexity of the smooth function, $c_{p}$ are constant coefficients to be estimated, and $b_{p}(\cdot)$ are the basis functions. Common basis functions include cubic regression splines (Durrleman and Simon,, 1989), thin-plate regression splines (Wood,, 2003), B-splines (De Boor and De Boor,, 1978), and penalised B-splines (“P-splines”) (Eilers and Marx,, 1996). The estimation of these spline functions, historically achieved via the backfitting algorithm (Hastie and Tibshirani,, 1986), can now be conducted through penalised likelihood maximisation (Wood,, 2011). This approach integrates a roughness penalty directly into the fitting criterion to automatically prevent overfitting. The theoretical and practical development of GAMs has been advanced substantially by Wood’s mgcv (Mixed GAM Computation Vehicle) (Wood,, 2017, 2012) package in the R programming language (R Core Team,, 2024).
@@ 2.2 Neural networks
Neural networks are a class of models inspired by the structure of the human brain, where information flows through the interaction of neurons (Goodfellow et al.,, 2016). Neural networks offer a flexible approach to modelling $f(\mathbf{x})$ by approximating complex, non-linear relationships through layers of interconnected nodes. A standard architecture is the feedforward multilayer perceptron (MLP) neural network. This architecture passes the data through a series of weighted transformations and non-linear activation functions. The default structure of a neural network is fully connected, essentially specifying all possible interactions (assuming a sufficiently complex neural network). An MLP with $L$ layers can be described by the equation:
$f(\mathbf{x})=h_{L}(\mathbf{b}_{L}+\mathbf{W}_{L}h_{L-1}(\dots\mathbf{b}_{2}+\mathbf{W}_{2}h_{1}(\mathbf{b}_{1}+\mathbf{W}_{1}\mathbf{x}))),$ (2.4)
where $L$ is the number of layers, $\mathbf{b}_{\ell}$ is a bias vector, $\mathbf{W}_{\ell}$ is a weight matrix, and $h_{\ell}(\cdot)$ is a non-linear activation function ($\ell=1,\ldots,L$). Figure 1 illustrates the architecture of a feedforward multilayer perceptron neural network. Neural networks are typically optimised via stochastic gradient descent and its variants (Bishop,, 1995; Bottou,, 2010), which iteratively update parameters to minimise a loss function using small, random subsets of data (batches).
@@
[…]
Figure 1: Feedforward multilayer perceptron neural network.
@@
RQ3 What are the potential drivers for the difference in performance of GAMs and NNs? Identify factors such as sample size, number of predictors, and dataset characteristics that influence model performance and examine how the relative performance of these models has evolved over time.
3 Research questions
Prior to carrying out a systematic review of comparisons of GAMs and neural networks, we defined key research questions to structure this review and facilitate a targeted comparison between the two model classes. Table 1 highlights our three central research questions along with the motivation behind each.
@@ 4 Methods
This review was reported following the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines (Page et al.,, 2021).
@@ 4.4 Selection process
To determine whether a document would be included in the final systematic review, each study was evaluated against the set of predefined eligibility criteria (Section 4.1) to ensure relevance. In order to assess whether a document satisfied these conditions, the entire document was reviewed. The first author conducted the screening of each paper to determine its eligibility for inclusion. In cases where inclusion was uncertain, the paper was discussed with the second and third authors, and a decision was made. This resulted in a total of 143 papers, from which comparisons on 430 datasets were included in our final systematic review (as many papers conduct comparisons on more than one dataset).
@@ 4.5 Data collection process
Data type Real or simulated data Manual Dataset
@@ 5 Study selection
Papers identified through database searching ($n=445$) Papers after duplicates removed ($n=444$) Full-text papers assessed for eligibility ($n=444$) Papers included in synthesis ($n=143$) Datasets included in synthesis ($n=429$) Identificationg Screening Includedg Duplicates removed ($n=1$) Full-text papers excluded ($n=301$): Papers not in English ($n=16$) Papers without empirical analysis between GAMs and NNs on same data ($n=235$) Papers lacking numeric values of performance metrics and/or author-reported statement on performance ($n=17$) Papers where results are in-sample ($n=9$) Papers where results are from simulated datasets only ($n=12$) Papers where full text was unavailable ($n=12$)
@@ 5.1 Quality assessment
The level of missingness in the literature raises concerns about the level of transparency of reporting in the current literature comparing GAMs to neural networks. Key methodological details, such as dataset characteristics and model complexity, are essential for evaluating the observed performance differences and for the reproducibility of the analyses carried out within these papers. The lack of standardised reporting also suggests a need for clearer guidelines and policies to promote complete reporting of experimental details in publications.
@@ 6 Study characteristics
Table 4 lists all 143 papers included in this systematic review, grouped by domain area.

### v2

@@ 1 Introduction
Over the past two decades, rapid advancements in machine learning and artificial intelligence have reshaped the abilities of predictive modelling. Among the most transformative of these developments are neural networks (RosenblattF1958TpAp; rumelhart1986learning), which have demonstrated high predictive performance across a wide range of tasks, including image recognition, natural language processing, and forecasting (abiodun2018state; liu2017survey). More recently, deep and alternative neural network architectures have been successful in a wide variety of tasks (arik2021tabnet; hollmann2022tabpfn; zhou2024deep; jia2024performance; hollmann2025accurate; CNN_tab_AnH_ILDO). Notwithstanding these advances, Multilayer Perceptrons (MLPs) remain the most common neural network baseline for tabular datasets (gorishniy2021revisiting), and this is also confirmed within the large scale systematic review of the current paper.
@@
In many cases, the primary objective when employing neural networks is to maximise predictive accuracy, with less emphasis placed on model interpretability. These models are often regarded as “black-box” algorithms: although their interconnected and non-linear structure allows for high flexibility, it also makes for a model that is difficult to interpret. One of the key strengths of neural networks is that, even with a single hidden layer, they have been proven to be universal approximators (cybenko1989approximation; hornik1989multilayer). This property implies that neural networks can approximate any continuous, single-valued function, and hence could be strong predictors. However, these theories of universality may require the neural network size to grow to infinity. Subsequent studies have shown that deep neural networks of a fixed size, i.e., neural networks with more than one hidden layer, can also achieve universal approximation, often with a drastically reduced number of parameters (lin2017does; mhaskar2016deep), thereby improving computational efficiency. Moreover, choromanska2015loss show that the loss landscape of deeper networks contains many local minima, which are not difficult to find using stochastic gradient descent, and perform similarly to the global minimum. Interestingly, for regression tasks, yun2019small prove that three-layer neural networks with ReLU (Rectified Linear Unit) activations can fit any arbitrary dataset when the number of hidden neurons equals the square root of the sample size, while schmidt2020nonparametric establish that sparsely connected deep ReLU networks can achieve optimal minimax estimation rate.
@@
The majority of neural network research has been conducted outside of the field of statistics, resulting in a limited incorporation of statistically-based methodologies and a disparity in the terminology used (hooker2021bridging). Neural networks, particularly MLPs, have emerged as strong alternatives to traditional statistical models, albeit they are often viewed as quite separate entities. However, a number of researchers have illustrated the similarities between neural networks and statistical models. NN_as_Statmod_WarrenS conducted a comprehensive comparison of various neural network models with traditional statistical approaches, showing that MLPs can be viewed as generalisations of non-linear regression and discriminant models. warner1996understanding highlight the parallels between neural networks and statistical regression models in terms of notation and implementation, showing that neural networks act as a type of nonparametric regression model. cheng1994neural inform statistical readers about neural networks, highlighting similarities with statistical methodology. More recently, ha2024fitting utilise deep neural networks within both a generalised non-linear modelling framework (extending classical generalised linear models) and a non-linear survival model (extending the Cox model). Furthermore, McInerney_2024 place neural networks within a statistical Gaussian likelihood-based framework, enabling input and architecture selection in a manner that is quite familiar within statistical modelling, but less so in neural networks research. Synthesizing these views, it becomes evident that neural networks are a generalisation of statistical approaches whereby they automatically learn complex non-linear effects and interactions within their hidden layers; this is in contrast to the manual specification of interactions required in traditional statistical modelling, which frequently utilises simpler linear effects for practicality.
@@
Given previous work comparing the mathematical structure of statistical models and neural networks, the question then arises as to how these methods compare with each other in practical applications. Indeed, several authors have made strides at answering this over the years. ANN_vs_Stats_medical conducted a review and analysis of the literature comparing neural networks with linear, logistic, and Cox regression in the context of medical research. STAT_Vs_NN_transport provided a survey of applications in the transportation area, comparing neural networks with linear or logistic regression predominantly. ALAKA2018164 carried out a systematic review of bankruptcy prediction studies, comparing multiple discriminant analysis and logistic regression with six artificial intelligence techniques and proposing a framework to guide tool selection and the use of hybrid models for improved predictive performance. shin2021machine carried out a systematic literature review of machine learning techniques (random forests, decision and regression trees, support vector machines, and neural networks) versus conventional statistical models (logistic, Cox, and Poisson regression) for predicting heart failure readmission and mortality. cho2021machine carried out a systematic review of the same aforementioned machine learning methods and conventional statistical models for the prediction of myocardial infarction readmission and mortality. sun2022comparing conducted a systematic review and meta-analysis of machine learning techniques, including random forests, support vector machines, neural networks, and decision trees, against statistical methods, linear and Cox regression, for predicting heart failure events. NAZARETH2023119640 conducted a systematic literature review examining machine and deep learning techniques including neural networks across six financial domains, comparing their effectiveness to traditional statistical approaches (linear and logistic regression) in tasks such as stock prediction, portfolio management, and bankruptcy forecasting. Similarly, ELAZAB2024124780 reviewed the use of machine and deep learning models including linear and logistic regression for Alzheimer’s disease diagnosis. More generally, paliwal2009neural carried out a review of comparative studies across various application domains, evaluating multilayered feedforward neural networks against linear regression, logistic regression, and discriminant analysis. Their review critically assessed the literature based on several key criteria, including sample size, adherence to the assumptions of the techniques, the method by which results are validated, the metrics used for comparison, and whether statistically significant differences were observed in the outcomes.
@@
Despite the interest in comparing neural networks and statistical models, prior reviews have predominantly narrowed their view to particular subject areas, such as transportation and heart health, as mentioned above. Additionally, these reviews have focused on quite simple statistical models, such as linear regression, which inherently lack the capabilities to model non-linear effects that may be underlying in the data. Therefore, such comparisons are not fair representations of what is possible in more advanced non-linear statistical models. Specifically, Generalised Additive Models (GAMs) (GAM; hastie1990generalized) are the most popular extension to linear regression, replacing linear terms with arbitrary non-linear functions. The strength of GAMs is their ability to deal with non-linear relationships between predictor and response variables, while maintaining interpretability through their additive structure.
@@
To date, there has been no systematic review of the literature comparing GAMs with neural networks in a wide range of application areas, and, therefore, our current paper closes this gap. The purpose of the systematic review is to be a source of information on the types of studies that have carried out comparisons of these two model classes and the reported findings within these studies. To move beyond just a summary of these papers, we have further extracted and analysed the values of the performance metrics provided within the papers, along with characteristics that can potentially explain observed differences (e.g., application area, year of study, sample size, number of predictor variables, and neural network complexity). This dataset has been made available for further research at https://github.com/jessicadoohan/gam-vs-nn-review.
@@ 2 Theoretical background
Before providing the details of our systematic review, we first provide some brief technical details on the models under consideration. Traditional statistical models and neural networks share the fundamental objective of modelling the relationship between a set of covariates and a response variable. Let $Y\in\mathbb{R}$ denote the response variable and $X=(X_{1},X_{2},\dots,X_{k})^{\top}\in\mathbb{R}^{k}$ the vector of $k$ covariates, where we restrict attention to tabular data. The general regression framework can be written as
@@
where GAMs utilise smooth functions $g_{j}(x_{j})$ ($j=1,\ldots,k$) to model non-linear relationships and $\textbf{NN}(x_{1},x_{2},\dots,x_{k})$ is a fully connected neural network (NN).
@@ 2.1 Generalised additive models (GAMs)
GAMs extend linear regression by allowing non-linear relationships between the predictor variables $\mathbf{x}$, and the response $Y$. As seen in Equation 2.2, smooth functions $g(\cdot)$ replace the constant coefficients of linear regression, which increases flexibility. Smooth functions can be approximated as a linear combination of a set of known basis functions,
@@
where $P$ is the number of specified functions, ($b_{1}(\cdot),\ldots,b_{P}(\cdot)$), controlling the maximum complexity of the smooth function, and the $c_{p}$ terms are associated constants to be estimated from data. Common basis functions include cubic regression splines (durrleman1989flexible), thin-plate regression splines (wood2003thin), B-splines (de1978practical), and penalised B-splines (“P-splines”) (eilers1996flexible). The estimation of these spline functions, historically achieved via the backfitting algorithm (GAM), can now be conducted through penalised likelihood maximisation (wood2011fast). This approach integrates a roughness penalty directly into the fitting criterion to automatically prevent overfitting. The theoretical and practical development of GAMs has been advanced substantially by Wood’s mgcv (Mixed GAM Computation Vehicle) (wood2017generalized; wood2012mgcv) package in the R programming language (R_cite).
@@ 2.2 Neural networks
Neural networks are a class of models inspired by the structure of the human brain, where information flows through the interaction of neurons (goodfellow2016deep). Neural networks offer a flexible approach to modelling $f(\mathbf{x})$ by approximating complex, non-linear relationships through layers of interconnected nodes. A standard architecture is the MLP neural network. This architecture passes the data through a series of weighted transformations and non-linear activation functions. The default structure of a neural network is fully connected, essentially specifying all possible interactions (assuming a sufficiently complex neural network). An MLP with $L$ layers can be described by the equation:
$f(\mathbf{x})=h_{L}(\mathbf{b}_{L}+\mathbf{W}_{L}h_{L-1}(\dots h_{\ell}(\mathbf{b}_{\ell}+\mathbf{W}_{\ell}h_{\ell-1}(\dots\mathbf{b}_{2}+\mathbf{W}_{2}h_{1}(\mathbf{b}_{1}+\mathbf{W}_{1}\mathbf{x}))))),$ (2.4)
where $\mathbf{b}_{\ell}$ is a bias vector, $\mathbf{W}_{\ell}$ is a weight matrix, and $h_{\ell}(\cdot)$ is a non-linear activation function ($\ell=1,\ldots,L$). Figure 1 illustrates the architecture of a multilayer perceptron neural network. Neural networks are typically optimised via stochastic gradient descent and its variants (bishop1995neural; bottou2010large), which iteratively update parameters to minimise a loss function using small, random subsets of data (batches).
@@
Figure 1: Multilayer perceptron neural network.
3 Research questions
Prior to conducting our systematic review of comparisons of GAMs and neural networks, we defined key research questions to structure this review and facilitate a targeted comparison between the two model classes. Table 1 highlights our three central research questions along with the motivation behind each.
@@ 3 Research questions
RQ3 What variables are associated with differences in performance of GAMs and NNs? Identify factors such as sample size, number of predictors, and dataset characteristics that may be associated with model performance, and examine how the relative performance of these models has evolved over time.
@@ 4 Methods
This review was reported following the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines (PRISMA).
@@ 4.4 Selection process
To determine whether a document would be included in the final systematic review, each study was evaluated against the set of predefined eligibility criteria described in Section 4.1. Defining the inclusion criteria in advance provides consistency selecting each paper, maintains relevance of the selected papers to our research questions, reduces potential selection bias, and ensures reproducibility of our systematic review. To assess whether a given document satisfied these conditions, the entire document was reviewed. The first author conducted the screening of each paper to determine its eligibility for inclusion. In cases where inclusion was uncertain, the paper was discussed with the second and third authors, and a decision was made. Twelve papers could not be retrieved and were therefore excluded. Given the total of 444 papers considered in our review, this exclusion is quite minor ($3\%$). However, it may introduce a small non-response bias that should be kept in mind when interpreting the subsequent analysis. Applying our inclusion criteria to these 444 papers resulted in a total of 143 papers, from which comparisons on 430 datasets were included in our final systematic review (as many papers conduct comparisons on more than one dataset).
@@ 5.1 Quality assessment
The level of missingness in the literature raises concerns about the level of transparency of reporting in the current literature comparing GAMs to neural networks. Notably, this pattern is consistent with broader reporting practices observed across scientific fields. For example, various systematic reviews have found high rates of unreported sample sizes (41% in johnson2017rapid and 12% in arcaya2016research). In andaur2022completeness, a systematic review of clinical prediction model reporting, found many instances of missing information, e.g., sample size (76% missing), predictor variables included (86% missing), model specification (95% missing). Our data extraction process is therefore subject to the same limitations imposed by the original reporting quality observed within the literature.
Key methodological details, such as dataset characteristics and model complexity, are essential for evaluating the observed performance differences and for the reproducibility of the analyses carried out within these papers. The lack of standardised reporting also suggests a need for using clear guidelines and policies to promote complete reporting of experimental details in publications of which researchers have made strides at formalising (TriPod; mitchell2019model).
@@ 6 Study characteristics
In this section we address RQ1, which asks about the key characteristics of papers comparing GAMs and NNs. We summarise paper-level and dataset-level attributes extracted to characterise the scope and structure of the existing literature. Table 4 lists all 143 papers included in this systematic review, grouped by domain area.
Table 4: Papers included in the systematic review.
Domain area Papers Datasets
Environmental & Agricultural Sciences 66 199
Aguilos2024, Beaugrand2024, Bonsoms20241777, Fan2024, Fu2024, Gazis20242473, Jung2024, Lei2024, Mouta202447, Senaratna2024, Yang2024, Ghorbani20246096, Adamova2023, Boudreault2023, DelSanto2023, Li2023, Yan2023339, Benemann2022189, Shafizadeh-Moghadam20221080, Yadav2022194, Evans2021, Ma2021, Martín2021289, Mondal2021, Phelps2021850, Tian2021, Zhang2021, Dodangeh2020, Dorich20201186, Mandal2020, Yadav2020129, Mugo2020, Nguyen2019, Liu2018572, Pourghasemi2018177, Corona-Núñez201769, Gao2017347, Pôças2017177, Ouarda20161553, Shortridge20162611, Vogel2016, Durocher20151561, Li2015312, Lin20154088, Xu201414, Gutiérrez-Estrada201219, Queirolo2012293, Haywood20111174, Palialexis2011165, Aertsen20101119, Gutiérrez-Estrada20101451, Ogawa-Onishi20101728, Tisseuil2010279, Virkkala2010269, Gutiérrez-Estrada2009116, Lee2009367, López-Darias20081027, Araújo20051504, Luoto2005299, Megrey20051256, Wang2005201, Thuiller20031353, Brosse20021033, Brosse2000441, Brosse19991293, Shatar1999249
Engineering, Mathematics, & Computer Science 35 141
Chen202421429, Kang2024, Nascimento2024288, Olca2024, Pachauri2024, Shojaeian20245371, Berrisch2023, Koh2023, Ledmaoui20231004, Montero2023222, Razaq20233957, Zhang20231, Zhang2023, Duan20223238, Mestanza2022397, Pachauri202224566, Dudek2021, Fang20213503, Kamis2021, Matsumoto2021, Yang2021, Daniel2020, Park2020, Ilseven2019, Tsang20185804, Ilseven20171, Martín2016173, Francisco-Fernández2015541, Yang2015108, Fukuda20131, Aertsen2011929, Berg2007129, Schlink2006547, Morlini2001169, Faraway1998231
Health & Biomedical Sciences 12 25
Francisco2024, Esmaeili2023423, Xu2023, Moslehi2022, Shoko2022534, Taylor2019, Allard2017338, Cengiz2012282, Gregori2011277, Gregori2009777, Giachino20071657, Goldfarb-Rumyantzev2002252
Other 30 65
Pasha2024, Roldán2024, Szeląg2024, Amara-Ouali20231272, Balan20232278, Bolleddula2023, Cai20231432, Krämer2023365, Mathur2023289, Sanii202268, Zhou2022, Dwivedi2021, Fotiadis2021, Frasier2021, Li20211364, Rengarajan2021, Zhu2021, Anglart20208433, Zhu20201, Zihan2020, Wheeler2019, Baquero2018, Luan2018, Davoudi2017568, Nyhan2014305, Liu20121431, Ryser2009187, Marmion20082241, Heikkinen2007347, Lee2002237